You bought a stock one year ago for ​$50.52 per share and sold it today for ​$45.96 per share. It paid a ​$1.77 per share dividend today.

a. What was your realized​ return?

b. How much of the return came from dividend yield and how much came from capital​ gain?
Answer #1

a.Realized rate=(End value-Beginning value+Dividends)/Beginning value

=(45.96-50.52+1.77)/50.52=(5.52%)(Approx)(Negative).

b.Dividend yield=Dividend/Beginning value

=(1.77/50.52)=3.50%(Approx).

Capital gains yield=(End value-Beginning value)/Beginning value

=(45.96-50.52)/50.52=(9.03%)(Approx)(Negative).